# Meaningful Learning: Advancing Abstract Reasoning in Large Language Models via Generic Fact Guidance

### v1

@@ A.1 Prompt for Experiment I
A.2 Prmpt for Experiment II
<|im_start|> system You are Orca, an AI language model created by Microsoft. You are a cautious assistant. You carefully follow instructions. You are helpful and harmless and you follow ethical guidelines and promote positive behavior.<|im_end|> <|im_start|> user Question: {premise} Hypothesis1 or Hypothesis2? Hypothesis1: {hypothesis1} Hypothesis2: {hypothesis2} Do you think {fact} can be used to answer this question? You answer should follow the format like ‘‘Answer: Yes or No.’’<|im_end|> <|im_start|> assistant
@@ A.2 Prmpt for Experiment II
Appendix B Details for Generic Facts Filtering
We apply the following steps to filter generics fact from GenericsKB:
@@ Appendix B Details for Generic Facts Filtering
Appendix C Prompt for AbsR Construction
@@
{number}
@@
Appendix D K-example and R-example
D.1 K-example
<|im_start|> system You are a cautious assistant. You carefully follow instructions. You are given a question, a few options, and a rule, you should follow the given rule to answer the question. Your response should follow the format like ‘‘Explanation: ___Answer: ____’’<|im_end|> <|im_start|> user Fact: Cookie files are simple text files that can be viewed in Windows Notepad. Question: What type of application can be used to open and view cookie files on a Windows computer? Options: A) Microsoft Excel B) Adobe Photoshop C) Windows Notepad D) 3D Modeling Software<|im_end|> <|im_start|> assistant Explanation: Windows Notepad is a text editor that is capable of opening and displaying the contents of simple text files, which is the format of cookie files. Answer: C) Windows Notepad<|im_end|>
D.2 R-example
<|im_start|> system You are a cautious assistant. You carefully follow instructions. You are given a question together with a few options, you should give an explanation first then answer the question. Your response should follow the format like ‘‘Explanation: ___Answer: ____’’<|im_end|> <|im_start|> user Question: What type of application can be used to open and view cookie files on a Windows computer? Options: A) Microsoft Excel B) Adobe Photoshop C) Windows Notepad D) 3D Modeling Software<|im_end|> <|im_start|> assistant Explanation: Windows Notepad is a text editor that is capable of opening and displaying the contents of simple text files, which is the format of cookie files. Answer: C) Windows Notepad<|im_end|>
Appendix E Templates for Evaluation
E.1 LLaMA-2
The following are multiple choice questions (with answers) about abstract algebra. Find the degree for the given field extension Q(sqrt(2), sqrt(3), sqrt(18)) over Q. A. 0 B. 4 C. 2 D. 6 Answer: A
E.2 Orca-2 and MeanLearn
<|im_start|> system The following are multiple choice questions (with answers) about abstract algebra.<|im_end|> <|im_start|> user Question: Find the degree for the given field extension Q(sqrt(2), sqrt(3), sqrt(18)) over Q. Options: A. 0 B. 4 C. 2 D. 6<|im_end|> <|im_start|> assistant Answer: A<|im_end|>

### v2

@@ B.1.1 Orca-2
B.2 Prompt for experiment II: Orca-2
B.2.1 Orca-2
<|im_start|> system
You are Orca, an AI language model created by Microsoft. You are a cautious assistant. You carefully follow instructions. You are helpful and harmless and you follow ethical guidelines and promote positive behavior.<|im_end|>
<|im_start|> user
You are given a fact, do you know this fact? Just answer Yes or No, do not give any additional information.
Fact: {fact} <|im_end|>
<|im_start|> assistant
B.2.2 LLaMA-2 and GPT-3.5
System : You are a helpful assistant.
User: You are given a fact, do you know this fact? Just answer Yes or No, do not give any additional information.
Fact: {fact}
@@ B.2.1 Orca-2
Appendix C Details for generic facts filtering
We apply the following steps to filter generics facts from GenericsKB:
@@ Appendix C Details for generic facts filtering
Appendix D Prompt for AbsR construction
@@
You are an expert in creating questions, you should first offer a question together with some options based on the fact the user gives. Second, you should give an answer and an explanation guided by the given fact. You can propose questions in any area, including but not limited to history, law, medicine, math, science, computer science, psychology, AI, politics, economics, etc. Your response should follow the following format: “Question: ____ Options:_____ Answer: _____ Explanation: _____”. NOTE that the fact should be an implicit explanation for obtaining the true answer, which means the fact SHOULD NOT appear explicitly in the questions or the options. The explanations should be short. Please create {number} examples.
@@
Appendix E Details of human evaluation
We choose three annotators with good backgrounds in textual inference and event reasoning. The whole test set of AbsR is involved in the human evaluation (200 instances). For each annotator, we pay \$10 per hour, while in our country, the minimum wage is less than \$5 per hour. All annotators agreed to let us use their annotations. The evaluation agreements are 95%, 91%, and 90% for the question answering, generic fact supportance, and diversity tasks, respectively.
Appendix F Instructions of human evaluation
F.1 Performance
You are given some questions each with some options, all the questions are about commonsense and are generated by GPT-4. Please choose the most plausible option for each question. Just type your choice (such as (A)) in the Answer column.
F.2 Support rate
You are given some questions each with some options and a generic fact, all the questions are about commonsense and are generated by GPT-4. Please judge whether the generic can support to answer the question. Just type your choice (Yes or No) in the Answer column.
F.3 Diversity
[…]
Appendix G K-example and R-example
G.1 K-example
<|im_start|> system
You are a cautious assistant. You carefully follow instructions. You are given a question, a few options, and a rule, you should follow the given rule to answer the question. Your response should follow the format like ‘‘Explanation: ___Answer: ____’’<|im_end|>
<|im_start|> user
Fact: Cookie files are simple text files that can be viewed in Windows Notepad.
Question: What type of application can be used to open and view cookie files on a Windows computer?
Options:
A) Microsoft Excel B) Adobe Photoshop C) Windows Notepad D) 3D Modeling Software<|im_end|>
<|im_start|> assistant
Explanation: Windows Notepad is a text editor that is capable of opening and displaying the contents of simple text files, which is the format of cookie files.
Answer: C) Windows Notepad<|im_end|>
G.2 R-example
<|im_start|> system
You are a cautious assistant. You carefully follow instructions. You are given a question together with a few options, you should give an explanation first and then answer the question. Your response should follow the format like ‘‘Explanation: ___Answer: ____’’<|im_end|>
<|im_start|> user
Question: What type of application can be used to open and view cookie files on a Windows computer?
Options:
A) Microsoft Excel B) Adobe Photoshop C) Windows Notepad D) 3D Modeling Software<|im_end|>
<|im_start|> assistant
Explanation: Windows Notepad is a text editor that is capable of opening and displaying the contents of simple text files, which is the format of cookie files.
Answer: C) Windows Notepad<|im_end|>
Appendix H Templates for evaluation
H.1 LLaMA-2, LLaMA-3, and MeanLearn (8B)
The following are multiple-choice questions (with answers) about abstract algebra.
Find the degree for the given field extension Q(sqrt(2), sqrt(3), sqrt(18)) over Q.
A. 0 B. 4 C. 2 D. 6
Answer: A
H.2 Orca-2 and MeanLearn (7B and 13B)
<|im_start|> system
The following are multiple-choice questions (with answers) about abstract algebra.<|im_end|>
<|im_start|> user
Question: Find the degree for the given field extension Q(sqrt(2), sqrt(3), sqrt(18)) over Q.
Options:
A. 0 B. 4 C. 2 D. 6<|im_end|>
<|im_start|> assistant
Answer: A<|im_end|>
Appendix I Training details of RoBERTa-Large-based clusterer
We use the RoBERTa-Large [32] released by Meta. The batch size is 256. We use AdamW [26] to optimize the model with a learning rate of 1e-5 for 5 epochs. The computing device is one NVIDIA A100-SXM-64GB. The running time is about 12 minutes.